OpenStudy (anonymous):

Determine the intercepts and use symmetry to assist in sketching the graph for the equation: y=3-x^2

OpenStudy (mattfeury):

to determine the intercepts, you should solve when x = 0 (the y intercept) and when y = 0 (the x intercept)

OpenStudy (anonymous):

it's going to open downward with your vertex at the +3y

OpenStudy (anonymous):

For some reason this is confusing me.

OpenStudy (anonymous):

ok if the number that is getting sq. is smaller than 0 it opens downward if it's bigger than 0 it opens up. the 2nd number has a varible on it. it's goign to control the width of the parabla. the constant is going to contro the shift in the vertex.

OpenStudy (anonymous):

I actually am getting more confused as to what happens with the x^2. I understand x=0 and y=0 to find the intercepts; however I can't seem to write it properly. This question is just a sample for me to figure out before I actually do my homework assignments.

OpenStudy (anonymous):

In addition, graphing for symmetry is really confusing me.

OpenStudy (anonymous):

in this case you y intercept is going to be 3, but i don't have a pencile or paper handy to find your x intercept. as far as graphing for symmetry you got me on that one. A parabla is symetrical by definition. if you don't have a graphing calculator. I'ts a downward opening parabla(maximum) with the vertex on the positive 3 line of the y axis. the x intercepts close to are close to +,- 2 . I hope this helps a little.

OpenStudy (anonymous):

It does help a bit. I am just going to have to figure this out on my calculator too. I'm actually doing this for someone else, so I sure would like to get it right. lol. Thank you very much.

OpenStudy (anonymous):

Pencil and paper? y = 0 => x^2 = 3 => x = plus or minus sqrt(3) ... And for the symmetry - if an equation is in x^2 and y^2 ONLY then it is symmetric in both the x and y axis (and the sign of x/y does not matter).

OpenStudy (anonymous):

If it is in only x^2 / y^2, then it is symmetric in the y/x axis, respectively, for the same reason.

OpenStudy (anonymous):

Wow, you are so awesome. Thank you so very much.